OpenStudy (anonymous):

Graph x2 + y2 = 9. What are its lines of symmetry?

jimthompson5910 (jim_thompson5910):

hint: \[\Large x^2 + y^2 = r^2\]

OpenStudy (anonymous):

so is every liine through the center a line of symmetry?

jimthompson5910 (jim_thompson5910):

yes that is correct. Why is that? What figure does this equation graph?

OpenStudy (anonymous):

its either

jimthompson5910 (jim_thompson5910):

What is the center of \[\Large x^2 + y^2 = 9\]

jimthompson5910 (jim_thompson5910):

and what is the radius of \[\Large x^2 + y^2 = 9\]

OpenStudy (anonymous):

is it the big circle or the little circle?

OpenStudy (anonymous):

Look at the first hint he gave you. Look at that equation and then look at the 9.

jimthompson5910 (jim_thompson5910):

what is the radius of \[\Large x^2 + y^2 = 9\]

jimthompson5910 (jim_thompson5910):

keep in mind it is in the form \[\Large x^2 + y^2 = r^2\]

OpenStudy (anonymous):

so its the little circle?

jimthompson5910 (jim_thompson5910):

what radius did you get?

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Yes the radius is 3. You're right it is the little circle because if you look it's radius is 3. Good job =D